Joseph is twelve years younger than Mack. Kevin is 1year younger than Joseph. The sum of their ages is sixty-two. How old is joseph.
Found 2 solutions by checkley71, josmiceli:
Answer by checkley71(8403)

M-12=J & K=J-1 & M+J+K=62 OR M=J+12 & K=J-1 THUS
J+12+J+J-1=62 OR 3J=62+1-12 OR 3J=51 OR J=17 THEREFORE M-12=17 OR M=17+12 OR
M=29 & K=17-1 OR K=16
PROOF 17+29+16=62 OR 62=62

Answer by josmiceli(6784)

J = M - 12
or, M = J + 12
K = J - 1
J + M + K = 62
find K and M in terms of J
J + (J + 12) + (J - 1) = 62
3J + 11 = 62
3J = 51
J = 17
--------
Find K and M
M = J + 12
M = 17 + 12
M = 29
---------
K = J - 1
K = 17 - 1
K = 16
----------
Now check answers
J + M + K = 62
17 + 29 + 16 = 62
checks
M = J + 12
29 = 17 + 12
checks
K = J - 1
16 = 17 - 1
checks